Exercise 25: Let f : [a, b] → R be continuous and ϕ : [a, b] → R be integrable and
non-negative, i.e. ϕ ≥ 0. Then there exists c ∈ (a, b) such that
the integral of f(x)Ï•(x)dx from a to b = f(c) the integral of Ï•(x)dx from a to b.
1
Expert's answer
2020-11-19T17:59:20-0500

Let "\int_a^b \phi(x)=0." Since for all "x \in[a, b], \phi(x)\geq0," then it suffices that "\forall x \in[a, b];\phi(x)=0."


Then, "\int_a^bf(x)\phi(x)dx=\int_a^bf(x).0dx=f(k).0=0"

Now suppose "\int_a^b\phi(x)dx>0" since "f" is continuous on "[a, b]" Then "f" is integrable on "[a, b]". Then by Extreme value theorem, there exists "m, M" such that

"f(m)=min f(x)\\
f(M)=max f(x)\\
x \in[a, b]"

Them we have the inequality

"f(m)\leq f(x) \leq f(M) ...........(1)"


Multiplying (1) by "\phi(x)" we have


"f(m)\phi(x) \leq f(x) \phi(x) \leq f(M)\phi(x)"

Integrating both sides, we have;


"f(m)\int_a^b\phi(x)dx \leq f(x) \int_a^b \phi(x)dx \leq f(M)\int_a^b\phi(x)dx"

Dividing by "\int_a^b\phi(x)dx," we have;


"f(m) \leq \frac{\int_a^bf(x)\phi(x)dx}{\int_a^b\phi(x)dx} \leq f(M)"


Then by the intermediate value theorem there exists "c\in(a, b)" such that


"\frac{\int_a^bf(x)\phi(x)dx}{\int_a^b\phi(x)dx} =f(c)"


"\implies \int_a^bf(x)\phi(x)dx=f(c)\int_a^b\phi(x)dx."